OpenStudy (anonymous):

ln 1/sqrt e = -1/2 write in exponential form pls show steps ln 1/sqrt e = -1/2 write in exponential form pls show steps @Mathematics

OpenStudy (anonymous):

\[\ln 1/\sqrt{e} =-1/2\]

OpenStudy (agreene):

\[\ln \frac1{\sqrt e}=-\frac12\] \[\ln e^{-\frac12}=-\frac12\] \[\cancel \ln \cancel e^{-\frac12}=-\frac12\] \[-\frac12=-\frac12\] \[QED\]

OpenStudy (anonymous):

what is QED?

OpenStudy (agreene):

"quod erat demonstrandum"

OpenStudy (agreene):

Or, "Thus I have shown" Its a way of ending a proof.

OpenStudy (anonymous):

ahh. i thought it was a cool answer.

OpenStudy (agreene):

http://en.wikipedia.org/wiki/Q.E.D. :)

OpenStudy (anonymous):

so 1/sqrt e is e^-1/2? or conversly, 1/sqrt a is a^-1/n?

OpenStudy (agreene):

\[\frac1a ≡ a^{-1}\] and \[\sqrt{a} ≡ a^{1/2}\] therefore: \[\large\frac1{\sqrt a} = a^{-\frac12}\]

OpenStudy (anonymous):

nice. thanks for your help. I struggled the first month in algebra.. really paying for it now with exponents and radicals.. especially with rational exponents

OpenStudy (agreene):

Whereas: \[\huge\sqrt[n]a ≡a^{\frac1n}\] and thus: \[\huge\frac{1}{\sqrt[n]a}=a^{-\frac1n}\] and no problem :)

OpenStudy (anonymous):

Ill have to add these to my big retriceformula sheet that I am trying to memorize!

OpenStudy (agreene):

hehe, once you get the hang of how some of these things work (this one especially) it will make your life a lot easier, because you'll see much easier ways of doing things than if you thought you were forced to use a fraction (you can just take it to a negative power ;) )